# Self-similar sets and Lipschitz graphs

### v1

@@ 1. Introduction
Rectifiable sets have been widely studied in the last century, following the pioneering work of Besicovitch, Federer, and Marstrand [Bes28, Fed47, Mar54], among many others. They have been since characterized and utilized as a class of “nice” sets for many problems in analysis. Their “bad” counterpart, purely unrectifiable sets, have also been studied for their remarkable measure theoretic properties. Attempting to provide a complete reference list for such statements is beyond our scope; see for instance [Mat95, Fal86] for an exposition of such topics.
Here we focus on one-dimensional sets in the plane with finite and positive measure ($1$-sets); those sets $E\subseteq\mathbb{R}^{2}$ with $\dim(E)=1$ and $0<\mathcal{H}^{1}(E)<\infty$. Here $\dim(\cdot)$ denotes the Hausdorff dimension and $\mathcal{H}^{s}$ denotes the $s$-dimensional Hausdorff measure, see Definition 2.2. Let $\mathcal{I}$ denote the collection of all Lipschitz images in the plane. That is, we say that $I\in\mathcal{I}$ if there exists a Lipschitz function $g:\mathbb{R}\to\mathbb{R}^{2}$ such that $I=g(\mathbb{R})$. We say that a $1$-set $E$ is rectifiable if there exists a countable collection of Lipschitz images $\left\{I_{i}\right\}_{i=1}^{\infty}\subseteq\mathcal{I}$ for which
@@
The main goal of this paper is to quantify the difference between rectifiable and purely unrectifiable sets and their relationships with the families of sets discussed above. As we show in Section 2, the heart of the matter will be Question 2, as Questions 1 and 3 are somewhat less involved. In the case of lines, the answer depends on the specific geometry of the set; while for the case of Lipschitz images, we show that for many purely unrectifiable sets, the intersection with the set has dimension 1, of course with zero measure. This result for Lipschitz images is given in Proposition 2.8. Our answer to Question 2 regarding Lischitz graphs is given in Theorem 2.10.
An important class of purely unrectifiable sets arise as the attractors of dynamical systems known as iterated function systems (IFS). Many well known fractals are attractors of IFS which satisfy the open set condition. See [Sch94, Sch96] for more details about the necessity of the open set condition. In this paper, we focus on $1$-dimensional attractors of iterated function systems that satisfy the open set condition (see Section 3 for definitions). By assuming the open set condition, the sets that we work with necessarily have two fundamental properties: self-similarity and Ahlfors regularity, both of which are key factors in our constructions.
A natural question in the theory of IFS is whether their attractors can be parametrized by nice curves. In this direction, Hata [Hat85] showed that if the attractor of an IFS is connected, then it is the continuous image of $[0,1]$. More recently, Badger and Vellis [BV21] improved the aforementioned result by showing that if an attractor is connected, then it is the image of a Hölder curve. Shaw and Vellis [SV24] provide sufficient conditions for the attractor of an infinite IFS to be parametrized by a Hölder curve.
@@ 1.1. Organization of the article
[…]
In Section 5, we present two graph constructions for the $4$-corner Cantor set, $\mathcal{C}_{4}$. Note that the set $\mathcal{C}_{4}$ was used as an example of a set with positive $1$-measure and zero analytic capacity in [Den32, Gar70]. The first construction described in Section 5 is ad hoc, while the second one illustrates the general method that is used in Section 6.
@@ 1.2. Notation
For a set $E\subseteq\mathbb{R}^{d}$, we write $\text{int}(E)$ to denote it interior, $\overline{E}$ to denote its closure, $\operatorname{conv}(E)$ to denote its convex hull, and $\operatorname{diam}(E)$ to denote its diameter. Given $r>0$, $E(r)$ denotes the closed $r$-neighborhood of $E$; that is, $E(r)=\left\{y\in\mathbb{R}^{d}:\left|x-y\right|\leq r\text{ for some }x\in E\right\}$. We write $B(x,r)=\left\{y\in\mathbb{R}^{d}:\left|x-y\right|\leq r\right\}$ to denote the closed ball of radius $r$ centered at $x\in\mathbb{R}^{d}$. A cube $Q\subseteq\mathbb{R}$ has sidelength denoted by $\ell(Q)$. Given a cube $Q$ and a constant $c>0$, we use the notation $c\,Q$ to denote the cube with the same center at $Q$ and sidelength $\ell(c\,Q)=c\,\ell(Q)$. If $\left\{E_{j}\right\}$ is a disjoint collection of sets, then we may write $\bigsqcup E_{j}$ to emphasize that the union is disjoint. If $E$ is a finite set, then we write #$\#E$ to denote the number of elements in $E$. For a Lebesgue measurable set $E\subseteq\mathbb{R}^{d}$, $|E|$ will denote the Lebesgue measure of $E$. Let $P_{\theta}:\mathbb{R}^{2}\to\mathbb{R}$ denote the orthogonal projection onto a line of angle $\theta$. That is, $P_{\theta}(z)=x\cos\theta+y\sin\theta$ for any point $z=(x,y)\in\mathbb{R}^{2}$. We may also write $P_{x}$ and $P_{y}$ to denote the projections onto that $x$- and $y$-axes, respectively. For inequalities, we write $A\lesssim B$ if there exists a constant $c>0$ such that $A\leq cB$.
1.3. Acknowledgements
@@ 2. Intersections with Lines and Lipschitz Images
The $s$-dimensional Hausdorff measure of a measurable set $E\subseteq\mathbb{R}^{d}$ is defined as
@@ 2.1. Intersections of purely unrectifiable 1-sets with lines
There exists an $s$-set $E\subseteq\left[0,1\right]^{2}$ such that $s\in[1,2]$ and
$\sup\left\{\dim(E\cap L):L\in A(2,1)\right\}=0.$
@@
$A$ :=$:=\left\{\sum_{j=1}^{\infty}\frac{a_{j}}{2^{j}}~{}:~{}a_{j}\in\{0,1\},\,a_{j}=0\mbox{ for }m_{2k}\leq j<m_{2k+1}\mbox{ for all }k\right\}$
$B$ :=$:=\left\{\sum_{j=1}^{\infty}\frac{b_{j}}{2^{j}}~{}:~{}b_{j}\in\{0,1\},\,b_{j}=0\mbox{ for }m_{2k-1}\leq j<m_{2k}\mbox{ for all }k\right\}.$
@@
For $k\geq 4$, let $\mathcal{C}_{k}$ be a k-square Cantor set, defined as $\mathcal{C}_{k}=\bigcap_{n=0}^{\infty}C_{n}$, where $C_{0}=[0,1]^{2}$, and $C_{1}$ is the union of $k$ squares, each of side-length $\frac{1}{k}$, $4$ of which are placed in the corners of $C_{0}$, and the remaining $k-4$ are evenly spaced along the $x$-axis. Then $C_{n}$ is defined recursively to consist of $k^{n}$ squares of sidelength $k^{-n}$. Images of the first two iterations of $\mathcal{C}_{6}$ are provided in Figure 1. If $L$ is the horizontal line through the origin, then $\dim\left(\mathcal{C}_{k}\cap L\right)=\frac{\log\left(k-2\right)}{\log k}$.
@@ 2.2. Intersections of purely unrectifiable 1-sets with Lipschitz images
For any Ahlfors regular $1$-set $E\subseteq\mathbb{R}^{2}$, there exists $I\in\mathcal{I}$ such that $\dim(E\cap I)=1$.
@@
(6) #$\#\mathcal{F}_{N_{k}}\in\left[M\lambda^{k-1}2^{N_{k-1}},M2^{N_{k-1}}\right].$
@@
By the 5-separation condition on $\mathcal{F}_{N_{k}}$, this union is disjoint, so the bounds from (5) and the inductive hypothesis described by (6) show that
#$\#\mathcal{G}^{0}_{n}$ $=\sum_{R\in\mathcal{F}_{N_{k}}}\#\mathcal{E}_{n}((1+2^{1-k})\cdot R)\in\left[Mc_{1}\lambda^{k-1}2^{n-k},MC_{1}2^{n-k}\right].$
Now define $\mathcal{G}_{n}\subseteq\mathcal{G}^{0}_{n}$ to be a maximal subcollection of cubes with the 5-separation condition. Since $\mathcal{G}_{n}$ is maximal and the set is in the plane, #$\#\mathcal{G}_{n}\geq\frac{1}{10^{2}}\#\mathcal{G}_{n}^{0}$ and we see that
#$\#\mathcal{G}_{n}$ $\in\left[Mc_{1}10^{-2}\lambda^{k-1}2^{n-k},MC_{1}2^{n-k}\right].$
Next, we construct $\mathcal{F}_{N_{k+1}}\subseteq\mathcal{E}_{N_{k+1}}$. Define $\mathcal{F}^{0}_{N_{k+1}}$ by choosing one cube, $q\in\mathcal{E}_{N_{k+1}}(2Q)$, for each of the $Q\in\mathcal{G}_{N_{k+1}-1}$. Since each $Q\in\mathcal{E}_{N_{k+1}-1}$, then the lower bound in (5) implies that such a $q=q(Q)$ always exists. Because $\mathcal{G}_{N_{k+1}-1}$ has the $5$-separation condition, then $Q\mapsto q(Q)$ is injective and #$\#\mathcal{F}^{0}_{N_{k+1}}=\#\mathcal{G}_{N_{k+1}-1}$. Recalling that $C_{1}\geq 1$, we reduce this set to $\mathcal{F}_{N_{k+1}}$ by selecting cubes so that #$\#\mathcal{F}_{N_{k+1}}\leq\frac{1}{C_{1}}\#\mathcal{F}^{0}_{N_{k+1}}$ and then
#$\#\mathcal{F}_{N_{k+1}}\in\left[M\lambda^{k}2^{N_{k}},M2^{N_{k}}\right],$
@@
Therefore, with $F_{k}:=\bigsqcup_{R\in\mathcal{F}_{N_{k}}}3R$, it holds that $F_{k+1}\subseteq F_{k}$ and by the 5-separation condition, these unions are disjoint. We note that each $F_{k}$ can be realized as a union of dyadic cubes in $\mathcal{Q}_{N_{k}}$.
@@
That is, $\mathcal{F}_{n}$ is the smallest collection of cubes in $\mathcal{Q}_{n}$ that covers $F_{k+1}$. The observation in (8) shows that $\bigcup_{Q\in\mathcal{F}_{n}}Q\subseteq F_{k}=\bigsqcup_{R\in\mathcal{F}_{N_{k}}}3R$ which implies that $\bigsqcup_{Q\in\mathcal{F}_{n}}3Q\subseteq\bigcup_{R\in\mathcal{F}_{N_{k}}}5R$. Let $G_{k}=\bigsqcup_{R\in\mathcal{F}_{N_{k}}}5R$ and note that the union is disjoint by the $5$-separation condition. For each $R\in\mathcal{F}_{N_{k}}$, there exists $x\in E\cap R$, so Ahlfors regularity implies that $\mathcal{H}^{1}\left(E\cap 5R\right)\leq\mathcal{H}^{1}\left(E\cap B(x,3\sqrt{2}\cdot 2^{-N_{k}})\right)\leq 3b\sqrt{2}\cdot 2^{-N_{k}}$. Therefore,
@@
Define $\mathcal{F}_{n}^{\prime}=\bigcup_{q\in\mathcal{F}_{N_{k+1}}}\{Q\in\mathcal{Q}_{n}~{}:~{}q\subseteq Q\}\subseteq\mathcal{F}_{n}$. If $Q\in\mathcal{F}_{n}\setminus\mathcal{F}_{n}^{\prime}$, then $Q$ must be a neighbor to some $Q^{\prime}\in\mathcal{F}_{n}^{\prime}$. Since each cube has eight neighbors, then #$\#\mathcal{F}_{n}^{\prime}\geq\frac{1}{9}\#\mathcal{F}_{n}$. Since $\mathcal{F}_{N_{k+1}}\subseteq\mathcal{E}_{N_{k+1}}$, then for each $Q\in\mathcal{F}_{n}^{\prime}$, we have $\mathcal{H}^{1}\left(E\cap Q\right)\geq\frac{a}{9}2^{-N_{k+1}}>0$. Following the argument in the proof of Lemma 2.9, either $Q\in\mathcal{E}_{n}$, or $Q$ has a neighbor in $\mathcal{E}_{n}$, so $\mathcal{H}^{1}\left(E\cap 3Q\right)\geq\frac{a}{9}2^{-n}$ for every $Q\in\mathcal{F}^{\prime}_{n}$. Thus, because $G_{k}\supseteq\bigcup_{Q\in\mathcal{F}_{n}}3Q\supseteq\bigcup_{Q\in\mathcal{F}_{n}^{\prime}}3Q$, we see that
@@
$\mathcal{N}_{N_{k}}$ =$=\left\{Q\in\mathcal{Q}_{N_{k}}:F\cap 3Q\neq\varnothing\right\}\subseteq\left\{Q\in\mathcal{Q}_{N_{k}}:F_{k}\cap 3Q\neq\varnothing\right\}=\left\{Q\in\mathcal{Q}_{N_{k}}:Q\subseteq\bigsqcup_{R\in\mathcal{F}_{N_{k}}}5R\right\}.$
In particular, #$\#\mathcal{N}_{N_{k}}\leq 25\#\mathcal{F}_{N_{k}}$. On the other hand, if $n\in\left(N_{k},N_{k+1}\right)$ for some $k\in\mathbb{N}$, then because $F_{k+1}\subseteq\bigcup_{Q\in\mathcal{F}_{n}}Q$, we see that
@@
and then #$\#\mathcal{N}_{n}\leq 9\#\mathcal{F}_{n}$. Now we have
@@
Therefore, Theorem 2.7 implies that there exists $I\in\mathcal{I}$ such that $F\subseteq I$.
By construction, $\left\{F_{k}\right\}_{k=1}^{\infty}$ is nested and each $F_{k}$ is a disjoint union of #$\#\mathcal{F}_{N_{k}}$ cubes with diameter $3\sqrt{2}\cdot 2^{-N_{k}}$. Thus, we may use Proposition 3.16 to estimate the dimension of $F$. Since $N_{k}=N_{k-1}+k=\frac{k(k+1)}{2}$ and #$\#\mathcal{F}_{N_{k-1}}\geq M\lambda^{k-2}2^{N_{k-2}}$, then
@@
from which it follows from pigeonholing that $\mathcal{H}^{1}\left(E\cap q_{k}\right)>\frac{a}{9}\ell(q)$ for some $k$. That is, for every $q\in\mathcal{E}^{\prime}_{n}\left(Q\right)$, there exists a neighbor $q^{\prime}\in\mathcal{E}_{n}\left(\left(1+2^{1-(n-m)}\right)\cdot Q\right)$. Since $\mathcal{H}^{1}\left(E\cap Q\right)>0$, then we must have that #$\#\mathcal{E}_{n}(\left(1+2^{1-(n-m)}\right)\cdot Q)\geq 1$. Since each $q^{\prime}$ can be a neighbor to at most eight cubes, then #$\#\mathcal{E}^{\prime}_{n}(Q)\leq 8\#\mathcal{E}_{n}(\left(1+2^{1-(n-m)}\right)\cdot Q)$. It follows that
@@
#$\leq\#\mathcal{E}_{n}\left(\left(1+2^{1-(n-m)}\right)\cdot Q\right)\left(\frac{8a}{9}+\sqrt{2}b\right)2^{-n}$

### v2

@@ 1. Introduction
Rectifiable sets have been widely studied in the last century, following the pioneering work of Besicovitch, Federer, and Marstrand [Bes28, Fed47, Mar54], among many others. They have since been characterized and utilized as a class of “nice” sets for many problems in analysis. Their “bad” counterpart, purely unrectifiable sets, have also been studied for their remarkable measure-theoretic properties. Attempting to provide a complete reference list for such statements is beyond our scope; see, for instance, [Mat95, Fal86] for an exposition of such topics.
Here we focus on one-dimensional sets in the plane with finite and positive measure ($1$-sets); those sets $E\subseteq\mathbb{R}^{2}$ with $\dim(E)=1$ and $0<\mathcal{H}^{1}(E)<\infty$. Here $\dim(\cdot)$ denotes the Hausdorff dimension and $\mathcal{H}^{s}$ denotes the $s$-dimensional Hausdorff measure, see Definition 2.2. Let $\mathcal{I}$ denote the collection of all Lipschitz images in the plane. That is, we say $I\in\mathcal{I}$ if there exists a Lipschitz function $g:[0,1]\to\mathbb{R}^{2}$ such that $I=g([0,1])$. We say that a $1$-set $E$ is rectifiable if there exists a countable collection of Lipschitz images $\left\{I_{i}\right\}_{i=1}^{\infty}\subseteq\mathcal{I}$ for which
@@
The main goal of this paper is to quantify the difference between rectifiable and purely unrectifiable sets and their relationships with the families of sets discussed above. As we show in Section 2, the heart of the matter will be Question 2, as Questions 1 and 3 are somewhat less involved. In the case of lines, the answer depends on the specific geometry of the set; while for the case of Lipschitz images, we show that for many purely unrectifiable sets, the intersection with the set has dimension 1, of course, with zero measure. This result for Lipschitz images is given in Proposition 2.8. Our answer to Question 2 regarding Lischitz graphs is given in Theorem 2.10.
An important class of purely unrectifiable sets arises as the attractors of dynamical systems known as iterated function systems (IFS). Many well-known fractals are attractors of IFS which satisfy the open set condition. See [Sch94, Sch96] for more details about the necessity of the open set condition. In this paper, we focus on $1$-dimensional attractors of iterated function systems that satisfy the open set condition (see Section 3 for definitions). By assuming the open set condition, the sets that we work with necessarily have two fundamental properties: self-similarity and Ahlfors regularity, both of which are key factors in our constructions.
A natural question in the theory of IFS is whether their attractors can be parametrized by nice curves. In this direction, Hata [Hat85] showed that if the attractor of an IFS is connected, then it is the continuous image of $[0,1]$. More recently, Badger and Vellis [BV21] improved the aforementioned result by showing that if an attractor is connected, then it is the image of a Hölder curve. Shaw and Vellis [SV24] provide sufficient conditions for the attractor of an infinite IFS to be parametrized by a Hölder curve. In [BK24], similar problems in metric spaces are explored for both Hölder and Lipschitz maps.
@@ 1.1. Organization of the article
The organization of the paper closely follows the steps we took in approaching the problem: In Section 2, we first address Questions 1 and 3. Section 3 provides definitions and preliminary results for iterated function systems, which can be of interest on their own. In Section 4, we demonstrate a general Lipschitz graph construction that parametrizes the $\limsup$ set of a suitable family of well-separated nested compact sets. In the sections that follow, we provide answers to Question 2.
In Section 5, we present two graph constructions for the $4$-corner Cantor set, $\mathcal{C}_{4}$. The set, also known as Garnett-Ivanov Cantor set, appears in works of Ivanov, translated in [Iva84], and later in Garnett [Gar70], who both proved the set has zero analytic capacity (while having positive length). Analogous sets, with scaling larger than $\frac{1}{4}$, were first considered by Denjoy [Den32] who proved those have positive continuous analytic capacity and positive Newtonian capacity.
The first construction described in Section 5 is ad hoc, while the second one illustrates the general method that is used in Section 6.
@@ 1.2. Notation
For a set $E\subseteq\mathbb{R}^{d}$, we write $\text{int}(E)$ to denote it interior, $\overline{E}$ to denote its closure, $\operatorname{conv}(E)$ to denote its convex hull, and $\operatorname{diam}(E)$ to denote its diameter. Given $r>0$, $E(r)$ denotes the closed $r$-neighborhood of $E$; that is, $E(r)=\left\{y\in\mathbb{R}^{d}:\left|x-y\right|\leq r\text{ for some }x\in E\right\}$. We write $B(x,r)=\left\{y\in\mathbb{R}^{d}:\left|x-y\right|\leq r\right\}$ to denote the closed ball of radius $r$ centered at $x\in\mathbb{R}^{d}$. A cube $Q\subseteq\mathbb{R}$ has side length denoted by $\ell(Q)$. Given a cube $Q$ and a constant $c>0$, we use the notation cQc\,Q to denote the cube with the same center at $Q$ and side length $\ell(c\,Q)=c\,\ell(Q)$. If $E$ is a finite set, then we write #E\#E to denote the number of elements in $E$. For a Lebesgue measurable set $E\subseteq\mathbb{R}^{d}$, $|E|$ will denote the Lebesgue measure of $E$. Let $P_{\theta}:\mathbb{R}^{2}\to\mathbb{R}$ denote the orthogonal projection onto a line of angle $\theta$. That is, $P_{\theta}(z)=x\cos\theta+y\sin\theta$ for any point $z=(x,y)\in\mathbb{R}^{2}$. We may also write $P_{x}$ and $P_{y}$ to denote the projections onto that $x$- and $y$-axes, respectively. For inequalities, we write $A\lesssim B$ if there exists a constant $c>0$ such that $A\leq cB$.
1.3. Funding and acknowledgements
@@ 2. Intersections with Lines and Lipschitz Images
The $s$-dimensional Hausdorff measure of a set $E\subseteq\mathbb{R}^{d}$ is defined as
@@ 2.1. Intersections of purely unrectifiable 1-sets with lines
There exists a set $E\subseteq\left[0,1\right]^{2}$ such that $\dim(E)\geq 1$ and
$\sup\left\{\dim(E\cap L):L\in A(2,1)\right\}=0,$
where we recall that $A(2,1)$ denotes the collection of all lines in $\mathbb{R}^{2}$.
@@
$A$ $:=\left\{\sum_{j=1}^{\infty}\frac{a_{j}}{2^{j}}~:~a_{j}\in\{0,1\},\,a_{j}=0\mbox{ for }m_{2k}\leq j<m_{2k+1}\mbox{ for all }k\right\}$
$B$ $:=\left\{\sum_{j=1}^{\infty}\frac{b_{j}}{2^{j}}~:~b_{j}\in\{0,1\},\,b_{j}=0\mbox{ for }m_{2k-1}\leq j<m_{2k}\mbox{ for all }k\right\}.$
@@
For $k\geq 4$, let $\mathcal{C}_{k}$ be a k-square Cantor set, defined as $\mathcal{C}_{k}=\bigcap_{n=0}^{\infty}C_{n}$, where $C_{0}=[0,1]^{2}$, and $C_{1}$ is the union of $k$ squares, each of side-length $\frac{1}{k}$, 44 of which are placed in the corners of $C_{0}$, and the remaining k4k-4 are evenly spaced along the $x$-axis. Then $C_{n}$ is defined recursively to consist of $k^{n}$ squares of side length $k^{-n}$. Images of the first two iterations of $\mathcal{C}_{6}$ are provided in Figure 1. If $L$ is the horizontal line through the origin, then $\dim\left(\mathcal{C}_{k}\cap L\right)=\frac{\log\left(k-2\right)}{\log k}$.
@@ 2.2. Intersections of purely unrectifiable 1-sets with Lipschitz images
For any bounded Ahlfors regular $1$-set $E\subseteq\mathbb{R}^{2}$, there exists $I\in\mathcal{I}$ such that $\dim(E\cap I)=1$.
@@
(6) $\mathcal{F}_{N_{k}}\in\left[M\lambda^{k-1}2^{N_{k-1}},M2^{N_{k-1}}\right].$
@@
By the 5-separation condition on $\mathcal{F}_{N_{k}}$, this union is of disjoint sets, so the bounds from (5) and the inductive hypothesis described by (6) show that
$\#\mathcal{G}^{0}_{n}$ $=\sum_{R\in\mathcal{F}_{N_{k}}}\#\mathcal{E}_{n}((1+2^{1-k})\cdot R)\in\left[Mc_{1}\lambda^{k-1}2^{n-k},MC_{1}2^{n-k}\right].$
Now define $\mathcal{G}_{n}\subseteq\mathcal{G}^{0}_{n}$ to be a maximal subcollection of cubes with the 5-separation condition. Since $\mathcal{G}_{n}$ is maximal and the set is in the plane, $\mathcal{G}_{n}\geq\frac{1}{10^{2}}\#\mathcal{G}_{n}^{0}$ and we see that
$\#\mathcal{G}_{n}$ $\in\left[Mc_{1}10^{-2}\lambda^{k-1}2^{n-k},MC_{1}2^{n-k}\right].$
Next, we construct $\mathcal{F}_{N_{k+1}}\subseteq\mathcal{E}_{N_{k+1}}$. Define $\mathcal{F}^{0}_{N_{k+1}}$ by choosing one cube, $q\in\mathcal{E}_{N_{k+1}}(2Q)$, for each of the $Q\in\mathcal{G}_{N_{k+1}-1}$. Since each $Q\in\mathcal{E}_{N_{k+1}-1}$, then the lower bound in (5) implies that such a $q=q(Q)$ always exists. Because $\mathcal{G}_{N_{k+1}-1}$ has the $5$-separation condition, then $Q\mapsto q(Q)$ is injective and $\mathcal{F}^{0}_{N_{k+1}}=\#\mathcal{G}_{N_{k+1}-1}$. Recalling that $C_{1}\geq 1$, we reduce this set to $\mathcal{F}_{N_{k+1}}$ by selecting cubes so that $\mathcal{F}_{N_{k+1}}\leq\frac{1}{C_{1}}\#\mathcal{F}^{0}_{N_{k+1}}$ and then
$\#\mathcal{F}_{N_{k+1}}\in\left[M\lambda^{k}2^{N_{k}},M2^{N_{k}}\right],$
@@
Therefore, with $F_{k}:=\bigcup_{R\in\mathcal{F}_{N_{k}}}3R$, it holds that $F_{k+1}\subseteq F_{k}$ and by the 5-separation condition, these unions are of disjoint sets. We note that each $F_{k}$ can be realized as a union of dyadic cubes in $\mathcal{Q}_{N_{k}}$.
@@
That is, $\mathcal{F}_{n}$ is the smallest collection of cubes in $\mathcal{Q}_{n}$ that covers $F_{k+1}$. The observation in (8) shows that $\bigcup_{Q\in\mathcal{F}_{n}}Q\subseteq F_{k}=\bigcup_{R\in\mathcal{F}_{N_{k}}}3R$ which implies that $\bigcup_{Q\in\mathcal{F}_{n}}3Q\subseteq\bigcup_{R\in\mathcal{F}_{N_{k}}}5R$. Let $G_{k}=\bigcup_{R\in\mathcal{F}_{N_{k}}}5R$ and note that the union is of disjoint sets by the $5$-separation condition. For each $R\in\mathcal{F}_{N_{k}}$, there exists $x\in E\cap R$, so Ahlfors regularity implies that $\mathcal{H}^{1}\left(E\cap 5R\right)\leq\mathcal{H}^{1}\left(E\cap B(x,3\sqrt{2}\cdot 2^{-N_{k}})\right)\leq 3b\sqrt{2}\cdot 2^{-N_{k}}$. Therefore,
@@
Define $\mathcal{F}_{n}^{\prime}=\bigcup_{q\in\mathcal{F}_{N_{k+1}}}\{Q\in\mathcal{Q}_{n}~:~q\subseteq Q\}\subseteq\mathcal{F}_{n}$. If $Q\in\mathcal{F}_{n}\setminus\mathcal{F}_{n}^{\prime}$, then $Q$ must be a neighbor to some $Q^{\prime}\in\mathcal{F}_{n}^{\prime}$. Since each cube has eight neighbors, then $\mathcal{F}_{n}^{\prime}\geq\frac{1}{9}\#\mathcal{F}_{n}$. Since $\mathcal{F}_{N_{k+1}}\subseteq\mathcal{E}_{N_{k+1}}$, then for each $Q\in\mathcal{F}_{n}^{\prime}$, we have $\mathcal{H}^{1}\left(E\cap Q\right)\geq\frac{a}{9}2^{-N_{k+1}}>0$. Following the argument in the proof of Lemma 2.9, either $Q\in\mathcal{E}_{n}$, or $Q$ has a neighbor in $\mathcal{E}_{n}$, so $\mathcal{H}^{1}\left(E\cap 3Q\right)\geq\frac{a}{9}2^{-n}$ for every $Q\in\mathcal{F}^{\prime}_{n}$. Thus, because $G_{k}\supseteq\bigcup_{Q\in\mathcal{F}_{n}}3Q\supseteq\bigcup_{Q\in\mathcal{F}_{n}^{\prime}}3Q$, we see that
@@
$\mathcal{N}_{N_{k}}$ $=\left\{Q\in\mathcal{Q}_{N_{k}}:F\cap 3Q\neq\varnothing\right\}\subseteq\left\{Q\in\mathcal{Q}_{N_{k}}:F_{k}\cap 3Q\neq\varnothing\right\}=\left\{Q\in\mathcal{Q}_{N_{k}}:Q\subseteq\bigcup_{R\in\mathcal{F}_{N_{k}}}5R\right\}.$
In particular, $\mathcal{N}_{N_{k}}\leq 25\#\mathcal{F}_{N_{k}}$. On the other hand, if $n\in\left(N_{k},N_{k+1}\right)$ for some $k\in\mathbb{N}$, then because $F_{k+1}\subseteq\bigcup_{Q\in\mathcal{F}_{n}}Q$, we see that
@@
and then $\mathcal{N}_{n}\leq 9\#\mathcal{F}_{n}$. Now we have
@@
Therefore, since $F\subseteq E$ and $E$ is bounded, Theorem 2.7 implies that there exists $I\in\mathcal{I}$ such that $F\subseteq I$.
By construction, $\left\{F_{k}\right\}_{k=1}^{\infty}$ is nested and each $F_{k}$ is a union of $\mathcal{F}_{N_{k}}$ disjoint cubes with diameter $3\sqrt{2}\cdot 2^{-N_{k}}$. Thus, we may use Proposition 3.16 to estimate the dimension of $F$. Since $N_{k}=N_{k-1}+k=\frac{k(k+1)}{2}$ and $\mathcal{F}_{N_{k-1}}\geq M\lambda^{k-2}2^{N_{k-2}}$, then
@@
from which it follows from pigeonholing that $\mathcal{H}^{1}\left(E\cap q_{k}\right)>\frac{a}{9}\ell(q)$ for some $k$. That is, for every $q\in\mathcal{E}^{\prime}_{n}\left(Q\right)$, there exists a neighbor $q^{\prime}\in\mathcal{E}_{n}\left(\left(1+2^{1-(n-m)}\right)\cdot Q\right)$. Since $\mathcal{H}^{1}\left(E\cap Q\right)>0$, then we must have that $\mathcal{E}_{n}(\left(1+2^{1-(n-m)}\right)\cdot Q)\geq 1$. Since each $q^{\prime}$ can be a neighbor to at most eight cubes, then $\mathcal{E}^{\prime}_{n}(Q)\leq 8\#\mathcal{E}_{n}(\left(1+2^{1-(n-m)}\right)\cdot Q)$. It follows that
@@
$\leq\#\mathcal{E}_{n}\left(\left(1+2^{1-(n-m)}\right)\cdot Q\right)\left(\frac{8a}{9}+\sqrt{2}b\right)2^{-n}$